OpenStudy (kj4uts):

There is no solution to the equation csc x = -1. True or false

OpenStudy (anonymous):

\[\csc(x)=1/\sin(x)\] \[\csc(x)=1/\sin(x)=-1\] \[\sin(x)=-1\]

OpenStudy (anonymous):

false

OpenStudy (anonymous):

because i have solution

OpenStudy (kj4uts):

Thank you!